# Variational Bayesian Optimal Experimental Design with Normalizing Flows

### v1

@@ 4.3 Case 3: Cathodic electrophoretic deposition (e-coating)
The physical model adopted for the e-coating process, consisting of a PDE, is the baseline model from [59] and also summarized in D.1. The parameters of interest, $\boldsymbol{\theta}=\{C_{v},j_{\min},Q_{\min}\}$, are the volumetric coulombic efficiency (m3/C), minimum current (A/m2), and minimum charge (C/m2), respectively. Each parameter is endowed with a truncated normal prior, with detailed prior parameters provided in D.1. A measurement on the current (mA) can be obtained at time $t$ (s), modeled as
@@
Figure 8 presents the convergence of the upper and lower bound estimates from Eqn. 19 and 20, along with a reference EIG value, when optimizing over $\boldsymbol{\lambda}$ at design $j_{0}=5.0$ mA. The reference EIG is obtained from a high-quality NMC with the ‘reuse’ technique described in [8] in order to accommodate the relatively high computational cost for each forward model evaluation (i.e., a PDE solve). Thus, even with $N_{\text{out}}=N_{\text{in}}=10^{5}$, the reuse technique only requires $10^{5}$ total forward model evaluations (instead of $10^{10}$, which would be impractical). We further provide comparison of the bounds between low-quality optimization that uses $N_{\text{opt}}=10000$ versus high-quality estimation that uses $N_{\text{opt}}=80000$ (the evaluation of the bounds are always performed with $N=10000$). Results indicate that increasing $N_{\text{opt}}$ tightens both bounds. Under the same $N_{\text{opt}}$, the upper bound appears to be tighter than the lower bound. To illustrate the ability of $q(\mathbf{y};\boldsymbol{\lambda})$ in representing the marginal likelihood $p(\mathbf{y}|\mathbf{d})$, their marginal distributions are shown in Fig. 9, where we see excellent agreement between the approximation and the true distributions obtained from direct Monte Carlo.
@@ 4.4 Case 4: aphid population
The OED problem is then to determine the optimal $k$ measurement times, $\mathbf{d}=\{t_{1},t_{2},...,t_{k}\}$, $t_{k}\in[0,50]$. Following [15], we solve the design optimization for $k=1$ and 2 (i.e., designing for respectively 1 and 2 measurement times) through grid search, while for $k=3$ and 4 we use the simultaneous perturbation stochastic approximation (SPSA) algorithm [63]. Table 2 presents the design results obtained using vOED-NFs and the LB-KLD method that was used in [15]. LB-KLD constructs another lower bound to EIG based on the entropy power inequality and entropy’s concavity property, and adopts the nearest neighbor based entropy estimator within the lower bound estimator. Note that LB-KLD does not involve tightening the lower bound and therefore is not a variational lower bound, but instead directly estimates the lower bound, here using $30000$ samples. For vOED-NFs, we set $N_{\text{opt}}=20000$ and $N=10000$ such that the total number of samples is the same. Implementation details can be found in E.1.
@@
Table 2: Case 4. The third column shows the optimal design $\mathbf{d}^{*}$ found by LB-KLD and vOED-NFs. The fourth column shows the mean of 10 $\widehat{U}_{\text{LB-KLD}}$ estimates with different random seeds (each with $30000$ samples for Monte Carlo estimation) at the optimal designs, and ($\cdot$) indicates the standard deviation originates from different random partitions of the same 30000 samples used for entropy estimation. The last column shows the mean of 10 $\widehat{U}_{L}$ estimates (each with $N_{\text{opt}}=20000$ and $N=10000$) at the optimal designs but different $\boldsymbol{\lambda^{*}}$. Here ($\cdot$) indicates the standard deviation due to 10 random seeds for different initialization when optimizing $\boldsymbol{\lambda^{*}}$, and evaluation is conducted on the same $N=10000$ samples.
@@ 5 Conclusions
We validated vOED-NFs against established methods in two benchmark problems, and demonstrated vOED-NFs on a PDE-governed application of cathodic electrophoretic deposition and stochastic modeling of aphid population with implicit likelihood. The findings suggested that 4–5 compositions of the coupling layers were adequate to achieve a lower bias compared to previous approaches. Furthermore, we illustrated that vOED-NFs produced approximate posteriors that matched very well with the true posteriors, able to capture non-Gaussian and multi-modal features effectively.
A limitation of vOED-NFs is the lack of analytical results connecting bound estimator quality to design optimization. For example, even if the lower bounds are not tight but the bound gap is consistent across $\mathscr{D}$, the design maximizer may still be similar to the true optimal design. Another limitation is that, to retain good accuracy under high-dimensional $\boldsymbol{\theta}$, complex NFs architectures with high-dimensional $\boldsymbol{\lambda}$ may be needed, making the optimization problem challenging. Future research exploring efficient and scalable transport map architectures in vOED, such as with rectification operator [36] and probability flow ODEs [65], will be useful. The adoption of vOED-NFs into other OED structures, such as goal-oriented OED [66] and sequential OED [67], will also be important.
@@ Acknowledgement
This work was supported at the University of Michigan by Ford Motor Company under the grant “Hybrid Physics-Machine Learning Models for Electrodeposition”.
@@ Appendix A Proof for Proposition 1
Appendix B Case 1
B.1 Hyperparameters
@@
B.2 EIG estimated on training sample
Figure 14 shows the EIG evaluated using the same samples it used for optimization (i.e., it is ‘testing’ on the same samples that it used for ‘training’). We see that in this case, smaller $N_{\text{opt}}$ can actually yield higher EIG estimates than Hi-NMC that uses $N_{\text{out}}=N_{\text{in}}=20000$ samples, indicating an overfitting phenomenon where $\widehat{U}_{L}(d;\boldsymbol{\lambda}^{*})$ could overestimate the EIG. Though in this case it does alter the location of the optimal design, the observation suggests the importance of evaluating EIG on a separate sample set than those used for optimization in order to reduce the bias in estimating the lower bound, especially when $N_{\text{opt}}$ is small.
@@ B.2 EIG estimated on training sample
Appendix C Case 2
C.1 Hyperparameters
@@
$N_{\text{opt}}$ $50000$
$N$ $10000$
$N_{\text{batch}}$ $2048$
@@
Appendix D Case 3
D.1 Forward model
@@
D.2 Hyperparameters for the 10-dimensional case
@@
D.3 Hyperparameters for the 50-dimensional case
@@
D.4 Posteriors from vOED-NFs with LSTM summary network and $N_{\text{opt}}=80000$
Figure 15 plots vOED-NFs posteriors when using LSTM summary network and $N_{\text{opt}}=80000$. The posteriors match the SMC posteriors (reference) better compared to those trained with smaller $N_{\text{opt}}$ and when $\mathbf{y}$ is higher dimensional (i.e., from Fig. 12).
@@ D.4 Posteriors from vOED-NFs with LSTM summary network and $N_{\text{opt}}=80000$
Appendix E Case 4
E.1 Hyperparameters
The hyperparameters for Sec. 4.4 are given in Table 11. Note LB-KLD [15] is computed using $10000$ prior samples, each associated with $3$ data samples such that the overall number of forward model runs is $30000$. For vOED-NFs, we keep the total number of forward model runs at $30000$ ($N_{\text{opt}}=20000$ and $N=10000$).

### v2

@@ 4.3 Case 3: Cathodic electrophoretic deposition (e-coating)
The physical model adopted for the e-coating process, consisting of a PDE, is the baseline model from [62] and also summarized in E.1. The parameters of interest, $\boldsymbol{\theta}=\{C_{v},j_{\min},Q_{\min}\}$, are the volumetric coulombic efficiency (m3/C), minimum current (A/m2), and minimum charge (C/m2), respectively. Each parameter is endowed with a truncated normal prior, with detailed prior parameters provided in E.1. A measurement on the current (mA) can be obtained at time $t$ (s), modeled as
@@
Figure 8 presents the convergence of the upper and lower bound estimates from Eqn. 19 and 20, along with a reference EIG value, when optimizing over $\boldsymbol{\lambda}$ at design $j_{0}=5.0$ mA. The reference EIG is obtained from a high-quality NMC with the ‘reuse’ technique described in [9] in order to accommodate the relatively high computational cost for each forward model evaluation (i.e., a PDE solve). Thus, even with $N_{\text{out}}=N_{\text{in}}=10^{5}$, the reuse technique only requires $10^{5}$ total forward model evaluations (instead of $10^{10}$, which would be impractical). We further provide comparison of the bounds between low-quality optimization that uses $N_{\text{opt}}=10^{4}$ versus high-quality estimation that uses $N_{\text{opt}}=8\times 10^{4}$ (the evaluation of the bounds are always performed with $N=10^{4}$). Results indicate that increasing $N_{\text{opt}}$ tightens both bounds. Under the same $N_{\text{opt}}$, the upper bound appears to be tighter than the lower bound. To illustrate the ability of $q(\mathbf{y};\boldsymbol{\lambda})$ in representing the marginal likelihood $p(\mathbf{y}|\mathbf{d})$, their marginal distributions are shown in Fig. 9, where we see excellent agreement between the approximation and the true distributions obtained from direct Monte Carlo.
@@ 4.4 Case 4: aphid population
The OED problem is then to determine the optimal $k$ measurement times, $\mathbf{d}=\{t_{1},t_{2},...,t_{k}\}$, $t_{k}\in[0,50]$. Following [16], we solve the design optimization for $k=1$ and 2 (i.e., designing for respectively 1 and 2 measurement times) through grid search, while for $k=3$ and 4 we use the simultaneous perturbation stochastic approximation (SPSA) algorithm [66]. Table 2 presents the design results obtained using vOED-NFs and the LB-KLD method that was used in [16]. LB-KLD constructs another lower bound to EIG based on the entropy power inequality and entropy’s concavity property, and adopts the nearest neighbor based entropy estimator within the lower bound estimator. Note that LB-KLD does not involve tightening the lower bound and therefore is not a variational lower bound, but instead directly estimates the lower bound, here using $3\times 10^{4}$ samples. For vOED-NFs, we set $N_{\text{opt}}=2\times 10^{4}$ and $N=10^{4}$ such that the total number of samples is the same. Implementation details can be found in F.1.
@@
Table 2: Case 4. The third column shows the optimal design $\mathbf{d}^{*}$ found by LB-KLD and vOED-NFs. The fourth column shows the mean of 10 $\widehat{U}_{\text{LB-KLD}}$ estimates with different random seeds (each with $3\times 10^{4}$ samples for Monte Carlo estimation) at the optimal designs, and ($\cdot$) indicates the standard deviation originates from different random partitions of the same 30000 samples used for entropy estimation. The last column shows the mean of 10 $\widehat{U}_{L}$ estimates (each with $N_{\text{opt}}=2\times 10^{4}$ and $N=10^{4}$) at the optimal designs but different $\boldsymbol{\lambda^{*}}$. Here ($\cdot$) indicates the standard deviation due to 10 random seeds for different initialization when optimizing $\boldsymbol{\lambda^{*}}$, and evaluation is conducted on the same $N=10^{4}$ samples.
@@ 5 Conclusions
We validated vOED-NFs against established methods in two benchmark problems, and demonstrated vOED-NFs on a PDE-governed application of cathodic electrophoretic deposition and stochastic modeling of aphid population with an implicit likelihood. The findings suggested that 4–5 compositions of the coupling layers were adequate to achieve a lower bias compared to previous approaches. Furthermore, we illustrated that vOED-NFs produced approximate posteriors that matched very well with the true posteriors, able to capture non-Gaussian and multi-modal features effectively.
A limitation of vOED-NFs is the lack of analytical results connecting bound estimator quality to design optimization. For example, even if the lower bounds are not tight but the bound gap is consistent across $\mathscr{D}$, the design maximizer may still be similar to the true optimal design. Another limitation is that, to retain good accuracy under high-dimensional $\boldsymbol{\theta}$, complex NFs architectures with high-dimensional $\boldsymbol{\lambda}$ may be needed, making the optimization problem challenging. An associated potential difficulty arises from the general susceptibility of neural network training to local minima. Furthermore, although we have illustrated greater sampling efficiency of vOED-NFs compared to NMC, it still require a non-trivial number of forward model runs to tighten the lower bound and achieve high-quality posteriors; further improving sampling efficiency will be important. Future research exploring efficient and scalable transport map architectures in vOED, such as with rectification operator [37] and probability flow ODEs [68], will be useful. The adoption of vOED-NFs into other OED structures, such as goal-oriented OED [69] and sequential OED [70], will also be important.
@@ Acknowledgement
This work was supported at the University of Michigan by Ford Motor Company under the grant “Hybrid Physics-Machine Learning Models for Electrodeposition”. This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of Michigan, Ann Arbor.
@@ Appendix A Proof for Proposition 1
Appendix B Additional Discussions for Invertible Neural Networks
By incorporating the data $\mathbf{y}$ into the coupling layer, the vOED-NFs can then produce a posterior distribution conditioned on that data. Correspondingly, Eqn. 27 and 28 become
$f_{1}(\widetilde{\boldsymbol{\Theta}}|\mathbf{y})$ $=\begin{bmatrix}\widetilde{\boldsymbol{\Theta}}_{1}\\ \mathbf{Z}_{2}=\widetilde{\boldsymbol{\Theta}}_{2}\odot\exp(s_{1}(\widetilde{\boldsymbol{\Theta}}_{1},\mathbf{y}))+t_{1}(\widetilde{\boldsymbol{\Theta}}_{1},\mathbf{y})\end{bmatrix},$ (39)
$f_{2}(f_{1}(\widetilde{\boldsymbol{\Theta}}|\mathbf{y}))$ $=\begin{bmatrix}\mathbf{Z}_{1}=\widetilde{\boldsymbol{\Theta}}_{1}\odot\exp(s_{2}(\mathbf{Z}_{2}),\mathbf{y})+t_{2}(\mathbf{Z}_{2},\mathbf{y})\\ \mathbf{Z}_{2}\end{bmatrix}.$ (40)
The invertibility is ensured as follows. Once a sample for $\widetilde{\boldsymbol{\Theta}}$ is obtained, the outputs from $s_{1}(\cdot),s_{2}(\cdot)$ and $t_{1}(\cdot),t_{2}(\cdot)$ are then fixed. The resulting transformation for $\mathbf{Z}_{2}=\widetilde{\boldsymbol{\Theta}}_{2}\odot\exp(s_{1}(\widetilde{\boldsymbol{\Theta}}_{1},\mathbf{y}))+t_{1}(\widetilde{\boldsymbol{\Theta}}_{1},\mathbf{y})$ is then a linear function with respect to $\widetilde{\boldsymbol{\Theta}}_{2}$ and therefore invertible.
Appendix C Case 1
C.1 Hyperparameters
@@
C.2 EIG estimated on training sample
Figure 14 shows the EIG evaluated using the same samples it used for optimization (i.e., it is ‘testing’ on the same samples that it used for ‘training’). We see that in this case, smaller $N_{\text{opt}}$ can actually yield higher EIG estimates than Hi-NMC that uses $N_{\text{out}}=N_{\text{in}}=2\times 10^{4}$ samples, indicating an overfitting phenomenon where $\widehat{U}_{L}(d;\boldsymbol{\lambda}^{*})$ could overestimate the EIG. Though in this case it does alter the location of the optimal design, the observation suggests the importance of evaluating EIG on a separate sample set than those used for optimization in order to reduce the bias in estimating the lower bound, especially when $N_{\text{opt}}$ is small.
@@ C.2 EIG estimated on training sample
Appendix D Case 2
D.1 Hyperparameters
@@
$N_{\text{opt}}$ $5\times 10^{4}$
$N$ $10^{4}$
$N_{\text{batch}}$ $2048$ (Table 1) / $1024$ (Fig. 7)
@@
Appendix E Case 3
E.1 Forward model
@@
E.2 Hyperparameters for the 10-dimensional case
@@
E.3 Hyperparameters for the 50-dimensional case
@@
E.4 Posteriors from vOED-NFs with LSTM summary network and $N_{\text{opt}}=8\times 10^{4}$
Figure 15 plots vOED-NFs posteriors when using LSTM summary network and $N_{\text{opt}}=8\times 10^{4}$. The posteriors match the SMC posteriors (reference) better compared to those trained with smaller $N_{\text{opt}}$ and when $\mathbf{y}$ is higher dimensional (i.e., from Fig. 12).
@@ E.4 Posteriors from vOED-NFs with LSTM summary network and $N_{\text{opt}}=8\times 10^{4}$
Appendix F Case 4
F.1 Hyperparameters
The hyperparameters for Sec. 4.4 are given in Table 11. Note LB-KLD [16] is computed using $10^{4}$ prior samples, each associated with $3$ data samples such that the overall number of forward model runs is $3\times 10^{4}$. For vOED-NFs, we keep the total number of forward model runs at $3\times 10^{4}$ ($N_{\text{opt}}=2\times 10^{4}$ and $N=10^{4}$).
@@ F.1 Hyperparameters
Appendix G Training Time for EIG Estimation
This section provides a sense of the computational cost in terms of “running time” for EIG evaluation at a specific design. We note that these figures are dependent on the implementation, such as which packages are being used, which training hyperparameters are adopted, whether code is optimized or parallelized, etc. Therefore, we suggest it to be used for awareness only, rather than direct comparisons. All numerical comparisons are run on the University of Michigan Great Lakes Slurm HPC Cluster nodes, with each node equipped with a single Nvidia Tesla A40 or V100 GPU.
For case 1 in Sec. 4.1, Fig. 2(a), the EIG evaluation time at a specific design for approaches (a)–(e) are approximately: (a) (Hi-NMC) 57 seconds, (b) (Lo-NMC) less than 0.1 seconds, (c) (Re-NMC) 16.5 seconds, (d) (vOED-G) 43 seconds, and (e) 1 minute 38 seconds. For this case, the forward model run is extremely fast and thus the computational cost from vOED is mainly contributed by the optimization process.
For case 2 in Sec. 4.2, Table 1, the improvement in EIG and posterior approximation of vOED-NFs over vOED-GG comes at the expense of increased training time, with vOED-GG requiring approximately 1 minute and 39 seconds per EIG evaluation, while vOED-NFs takes around 5 minutes and 56 seconds under the selected hyperparameters.
For case 3 in Sec. 4.3, Fig. 8, the training times for (a) vsOED-NFs (upper), $N_{\text{opt}}=10^{4}$, (b) vsOED-NFs (upper), $N_{\text{opt}}=8\times 10^{4}$, (c) vsOED-NFs (lower), $N_{\text{opt}}=10^{4}$, (d) vsOED-NFs (lower), $N_{\text{opt}}=8\times 10^{4}$ are respectively 2 minutes 32 seconds, 5 minutes 59 seconds, 3 minutes 50 seconds, and 6 minutes 48 seconds under the selected hyperparameters. Whereas for the Hi-NMC estimator which requires $10^{5}$ forward model runs in total, each run involves solving a PDE and takes approximately 1.5 seconds. Therefore, the additional time induced by training NFs is trivial compared to the overall computational cost of running the forward models for Hi-NMC.
For case 3, Fig. 12, with LSTM incorporated via summary network for evaluating the EIG lower bounds, the training times for (a) vOED-NFs, $N_{\text{opt}}=10^{4}$, (b) vOED-NFs, $N_{\text{opt}}=8\times 10^{4}$, (c) vOED-NFs-LSTM, $N_{\text{opt}}=10^{4}$, (d) vOED-NFs-LSTM, $N_{\text{opt}}=8\times 10^{4}$ are respectively 4 minutes 33 seconds, 6 minutes 19 seconds, 5 minutes 13 seconds, and 7 minutes 38 seconds under the selected hyperparameters.
Lastly, for case 4 in Sec. 4.4, Fig. 13, the training time for vOED-NFs is approximately 15 seconds.